As Gina was working in her new yard, she decided she needed a little more color. When she went to the garden center at Costco, she spotted 2 azalea, 1 rhododendron and 2 daphne that she liked. If an experiment consits of her choosing 2 of the above ploants,
a) What is the sample space of this experiment?
b) What is the probability of selecting two plants not of the same type?

Answer by stanbon(75887)

As Gina was working in her new yard, she decided she needed a little more color. When she went to the garden center at Costco, she spotted 2 azalea, 1 rhododendron and 2 daphne that she liked. If an experiment consists of her choosing 2 of the above ploants,
a) What is the sample space of this experiment?
# of ways to pick the 1st plant = 5
# of ways to pick the 2nd plant = 4
# of ways to pick the two plants 5*4 = 20
There are 20 pairs in the sample space.
----------------------------------
b) What is the probability of selecting two plants not of the same type?
Two of the 20 pairs have plants of the same type.
There are 18 pairs that have plants that are not of the same type.
Prob (selecting two not of the same type) = 18/20 = 9/10
=================
Cheers,
Stan H.